# Adaptive Reduced Multilevel Splitting

### v1

@@ 2.3 Snapshot sampling and learning function
$X_{k}\sim\frac{1}{\mu^{(k)}_{l^{(k)}}(\mathrm{e}^{\tau^{(k)}E^{(k)}})}\mathrm{e}^{\tau^{(k)}E^{(k)}}\mu^{(k)}_{l^{(k)}},$
with $\tau^{(k)}\geq 0$. The above penalization is similar to the so-called learning function in the literature on ’active learning’ [19]. In this work, we simply choose an exponential function of the error $E$ with a penalty parameter $\tau^{(k)}$ adaptively chosen at iteration $k$ of the algorithm. In particular, $\tau^{(k)}$ will be chosen to be high in the beginning of the algorithm and then be reduced to 00 at some point. The idea here is that $\tau^{(k)}$ enables to find a compromise between two objectives: first i) the snapshots are used to refine the reduced model $S^{(k)}$, and ii) the snapshots are also used in estimation by importance sampling. A large $\tau^{(k)}$ is associated with the objective i). Indeed, favoring candidates with high errors in a $\mu^{(k)}_{l^{(k)}}$-distributed sample enables a faster learning of the true score $S^{\ast}$. A small or vanishing $\tau^{(k)}$ is associated with objective ii). Indeed, the distribution $\mu^{(k)}_{l^{(k)}}$ is constructed to be a robust importance sampling distribution for the target $\eta_{l^{(k)}}^{\ast}$ so that states with smaller error are needed for estimation, ensuring a limited variance for importance sampling.
@@
2.4 Bridging the score approximations
We will now focus on what happens after the update of the reduced model that follows the sampling of a new snapshot (true model). We are given a certain level $l^{(k)}$ and an associated importance distribution $\mu_{l^{(k)}}^{(k)}$. The problem is to find a new initial level which we will denote $l_{\mathrm{bridge}}(k+1)$ and then sample the new importance distribution $\mu_{l_{\mathrm{bridge}}(k+1)}^{(k+1)}$ associated with the updated score approximation $S^{(k+1)}$. This initial level will then be used in the subsequent $k+1$ iteration to determine (with definition (2.2)) the critical level at iteration $k+1$: $l^{(k+1)}=l^{(k+1)}(e,l_{\mathrm{bridge}}(k+1))$ for $S^{(k+1)}$. This will complete the main loop of the algorithm.
However, if one tries to sample $\mu_{l_{\mathrm{bridge}}(k+1)}^{(k+1)}$ starting from $\mu_{l^{(k)}}^{(k)}$, using some form of importance or sequential sampling, there is an issue with inclusion. For instance, simply setting the initial level ${l_{\mathrm{bridge}}(k+1)}$ to $l^{(k)}$, the nesting $\{S^{(k+1)}>l_{(k)}\}\subseteq\{S^{(k)}>l^{(k)}\}$ will generally not be valid, making the task infeasible.
@@ 2.4 Bridging the score approximations
$\mathrm{Ent}\left(\mu_{l_{\mathrm{bridge}}(k+1)}^{(k+1)}\mid\mu_{l^{(k_{\mathrm{bridge}})}}^{(k_{\mathrm{bridge}})}\right)\leq-\ln(1-\theta).$ (2.5)
This condition ensures that a relatively straightforward importance sampling procedure enables to obtain the initial importance sampling distribution $\mu_{l_{\mathrm{bridge}(k+1)}}^{(k+1)}$ of the forthcoming iteration, from a previous importance distribution $\mu_{l^{(k_{\mathrm{bridge}})}}^{(k_{\mathrm{bridge}})}$. In particular, the nesting condition $\{S^{(k+1)}\geq{l_{\mathrm{bridge}}(k+1)}\}\subseteq\{S^{(k_{\mathrm{bridge}})}>l^{(k_{\mathrm{bridge}})}\}$ will hold true.
In addition of the condition (2.5), we want to also ensure that the two conditions in (2.2) used to define the maximal achievable level $l^{(k+1)}$ hold for this initial level ${l_{\mathrm{bridge}}(k+1)}$. This is necessary to ensure formally that ${l_{\mathrm{bridge}}(k+1)}\leq l^{(k+1)}$. Explicitly:
$\mathrm{Ent}(\check{\eta}^{(k+1)}_{l_{\mathrm{bridge}}(k)}\mid\mu^{(k+1)}_{l_{\mathrm{bridge}}(k+1)})\leq e,$ (2.6)
$\mathrm{Ent}(\hat{\eta}^{(k+1)}_{l_{\mathrm{max}}}\mid\mu^{(k+1)}_{l_{\mathrm{bridge}}(k+1)})<+\infty.$ (2.7)
A pair $(k_{\mathrm{bridge}},{l_{\mathrm{bridge}}(k+1)})\in\llbracket 0,k\rrbracket\times\mathbb{R}$ satisfying (2.5), (2.6) and (2.7) determine two consecutive proposal distributions $\mu_{l^{(k_{\mathrm{bridge}})}}^{(k_{\mathrm{bridge}})}$ and $\mu^{(k+1)}_{l_{\mathrm{bridge}}(k+1)}$ “bridging” the score approximations $S^{(k_{\mathrm{bridge}})}$ and $S^{(k+1)}$. This pair can be determined according to various strategies.
  • A trivial feasible pair is to set $k_{\mathrm{bridge}}=0$ such that $l^{(k_{\mathrm{bridge}})}=l_{\mathrm{bridge}}(k+1)=-\infty$. As explained in Section 4, this will account in practice to restart the AMS simulation with the new score $S^{(k+1)}$ from the beginning until it reaches the critical level $l^{(k+1)}=l^{(k+1)}(e,-\infty).$ However, this option can lead to a significant computational overhead if the complexity of the reduced score evaluation is not negligible.
  • A non trivial feasible pair of interest is given by the largest (lexicographic) pair $(k_{\mathrm{bridge}},{l_{\mathrm{bridge}}(k+1)})\in\llbracket 0,k\rrbracket\times\mathbb{R}$ satisfying the three conditions (2.5), (2.6) and (2.7). This means first selecting the importance sampling distribution $\mu_{l^{(k_{\mathrm{bridge}})}}^{(k_{\mathrm{bridge}})}$, among the already computed distributions, from the most recent to the oldest, and then to search the largest level ${l_{\mathrm{bridge}}(k+1)}$ such that the three conditions hold. We will see in Section 4 that in practice, this solution will lead to an important computation saving, but will require nevertheless to save the history of the previous AMS simulations.
@@ 2.5 Stopping criteria
In practical situations, a maximum level $l_{\mathrm{max}}$ is setting the range of the computation of interest: one is interested in computing the rare event probability $p^{\ast}_{l_{\mathrm{max}}}$. The level $l^{(k)}(e,l_{0})$ defined in (2.2) is constrained to stay below $l_{\mathrm{max}}$, and the definition of $l^{(k)}(e,l_{0})$ is modified as the minimum between $l_{\mathrm{max}}$ and the critical level given by (2.2).
The algorithm can thus continue as long as one wishes, for instance until enough snapshot samples are obtained.
In addition, in order to avoid unnecessarily enriching an already accurate reduced score, it is possible to stop updating the reduced score if the worst-case log cost at the critical level $l^{(k)}$ is almost zero and if this critical level is just below $l_{\max}$. More precisely we define the update stopping index222 We call a stopping index any exit criteria of the algorithm that is independent of the future snapshots. as the first index $k$ such that
@@
where $\epsilon>0$ is a given threshold parameter close to the machine precision. We will however see in Section 4 that using such a stopping of update of the reduced score, the normalization constant is no longer updated resulting in a sustained variance while new snapshots are sampled. This sustained variance is however unavoidable when using a non trivial bridging procedure. We will however see in Section 4 that using the latter stop-update rule with a non-trivial bridge usually largely compensate additional variance with computational (and memory) savings.
3 An Idealized Scheme
3.1 Iterative importance sampling algorithm
We discuss in this section the estimation of the probability $p^{\ast}_{l_{\mathrm{max}}}$. It is more convenient to work with the un-normalized version of the target distribution $\eta^{\ast}_{l_{\mathrm{max}}}$ defined in (1.1), which will be denoted from now on by $\gamma^{\ast}_{l_{\mathrm{max}}}$ and defined by
@@ 3.1 Iterative importance sampling algorithm
for a maximum level of interest ${l_{\mathrm{max}}}\in\mathbb{R}$ and some test function $\varphi$. Note that $p^{\ast}_{l_{\mathrm{max}}}=\gamma^{\ast}_{l_{\mathrm{max}}}(\,{\rm l}{\rm l})$.
Based on the concepts introduced in Section 2, we propose in this section an idealized algorithm that will lead to an estimation of $\gamma^{\ast}_{l_{\mathrm{max}}}$ based on importance sampling. The algorithm defines at each iteration $k\geq 1$ some random level $l^{(k)}\in]-\infty,l_{\mathrm{max}}]$. To this level is associated the proposal distribution $\mu^{(k)}_{l^{(k)}}$ as defined by (1.3) (the distribution defined by conditioning on reduced scores $S^{(k)}$ greater than $l^{(k)}$). $l^{(k)}$ will also be associated to the pessimistic surrogate $\check{\eta}_{l^{(k)}}^{(k)}$ defined by (2.3). The algorithm will also make use of the optimistic surrogate $\hat{\eta}_{l_{\mathrm{max}}}^{(k)}$ defined in (2.4).
At iteration $k$ of the algorithm, a snapshot $X_{k+1}\in\mathcal{X}$ is drawn according to the penalized importance distribution $\propto\mathrm{e}^{-\tau^{(k)}S^{(k)}}\mu_{l^{(k)}}^{(k)}$, where $\tau^{(k)}$ is some adaptive “learning” parameter. The algorithm will use for estimation the full history of snapshots.
The algorithm is presented below. It is an idealized concept and, in practice, it is not directly tractable. Indeed it assumes exact evaluation of expectations and exact sampling according to the proposal and the surrogate distributions which are defined up to a normalization constant. In Section 4, we will propose a practical implementation of this idealized algorithm using an AMS (adaptive Sequential Monte Carlo) methodology. The latter will simulate the flow of proposals for growing level parameters $l$ by providing Monte Carlo samples with approximately correct distributions for large sample sizes.
Algorithm 1 (Iterative importance sampling)
@@
1:level of interest $l_{\mathrm{max}}$, initial reduced score $S^{(1)}$ and error $E^{(1)}$ functions, budget of snapshots $K$, worst-case log cost threshold $e$, reduced score precision $\epsilon\ll e$, number of hits of rare event before estimation $J$, quantile $\theta=M/N\in(0,1)$ for the bridge between proposals $\triangleright$ ( initialization)
2:Set $l^{(0)}=l_{\mathrm{bridge}}(1)=-\infty$, $H=0$, $j=0$, $k=0$, $\tau=+\infty$, ${Z_{l}}^{(1)}=1$
@@
5:    Simulate the flow of proposal distributions $\mu^{(k)}_{l}$ for $l\leq l^{(k)}$. The maximal level $l^{(k)}$ must be a level $l\in[l_{\mathrm{bridge}}(k),l_{\mathrm{max}}]$ satisfying both of the following two conditions:
$\mathrm{Ent}(\check{\eta}^{(k)}_{l}\mid\mu^{(k)}_{l})\leq e$
@@
$l^{(k)}$ is defined as the smallest $l\in[l_{\mathrm{bridge}}(k),l_{\mathrm{max}}]$ (in practice $l^{(k)}=l-\delta$ for some $\delta>0$ very small) violating one the above two conditions; if search fails, then $l^{(k)}=l_{\mathrm{max}}$.
@@
15:    if $j=J$ then set the learning function parameter as $\tau=0$ end if $\triangleright$ ( importance sampling estimation, see Section 3.2)
16:    if  $j>J$ then
@@
$\hat{\gamma}^{H}_{l_{\max}}(\varphi)\leftarrow\frac{(H-1)\hat{\gamma}^{H-1}_{l_{\max}}(\varphi)+Z^{{(k)}}_{l^{(k)}}\varphi(X)\,{\rm l}{\rm l}_{S^{\ast}(X)>l_{\max}}}{H}$
@@
22:         Find the largest (lexicographic) pair $(k^{\prime},l^{\prime})\in\llbracket 0,k\rrbracket\times\mathbb{R}$ satisfying the following three conditions:
@@
$\mathrm{Ent}(\check{\eta}^{(k+1)}_{l^{\prime}}\mid\mu^{(k+1)}_{l^{\prime}})\leq e,$
@@
and denote the solution pair by $(k_{\mathrm{bridge}}(k+1),l_{\mathrm{bridge}}(k+1))\leftarrow(k^{\prime},l^{\prime}).$
23:         Simulate the proposal $\mu^{(k+1)}_{l_{\mathrm{bridge}}(k+1)}$ using $\mu^{(k_{\mathrm{bridge}}(k+1))}_{l^{k_{\mathrm{bridge}}(k+1)}}$; then compute the associated normalization $Z^{(k+1)}_{l_{\mathrm{bridge}}(k+1)}$.
24:    else
25:         Set $l_{\mathrm{bridge}}(k+1)=l^{(k)}$
26:         Set $Z^{(k+1)}=Z^{(k)}$ and $\mu^{(k+1)}_{l_{\mathrm{bridge}}(k+1)}=\mu^{(k)}_{l^{(k)}}$
27:    end if
28:end while
29:Return Estimate $\hat{\gamma}^{H}_{l_{\max}}(\varphi)$
3.2 The estimator
As defined by the algorithm in the previous section, the importance sampling estimator $\hat{\gamma}_{l_{\max}}^{H}$ of $\gamma^{\ast}_{l_{\max}}$ is defined using a random iteration index which we will denote by $K_{\mathrm{min}}$. Only the snapshots sampled at and after iteration $K_{\mathrm{min}}+1$ will be used to build our importance sampling estimator. More precisely, $K_{\mathrm{min}}$ must be of optional type, in the sense that $K_{\mathrm{min}}$ is defined as the first iteration index $k$ for which a certain property depending on the snapshots $X_{1},\ldots,X_{k}$ holds true.
We propose the following option: $K_{\mathrm{min}}$ is the first iteration index $k$ such that the score of the past snapshots have reached the level $l_{\max}$ (the rare event) $J$ times, for some prescribed $J\geq 1$. At iteration $K_{\mathrm{min}}+H$ (for some $H>0$) of the algorithm, the estimator of $\gamma^{\ast}_{l_{\mathrm{max}}}$ is then given by:
$\hat{\gamma}^{H}_{l_{\mathrm{max}}}\displaystyle\mathop{=}^{{\rm def}}\frac{1}{H}\sum_{k^{\prime}=K_{\mathrm{min}}+1}^{K_{\mathrm{min}}+H}f_{l_{\mathrm{max}}}^{(k^{\prime})}(X_{k^{\prime}+1})\delta_{X_{k^{\prime}+1}},$ (3.1)
where the IS density ratio is given by
$f_{l_{\mathrm{max}}}^{(k)}(x)\displaystyle\mathop{=}^{{\rm def}}\frac{\mu_{{l^{(k)}}}^{{(k)}}(\mathrm{e}^{\tau^{(k)}S^{(k)}})}{\mathrm{e}^{\tau^{(k)}S^{(k)}(x)}}\frac{\,{\rm l}{\rm l}_{S^{\ast}\geq{l_{\mathrm{max}}}}\mathrm{d}\pi}{d\mu_{l^{(k)}}^{(k)}}(x).$
Note that in the algorithm: i) we have ensured that $\eta_{l_{\mathrm{max}}}^{\ast}\ll\mu_{l^{(k)}}^{(k)}$, and ii) $f_{l_{\mathrm{max}}}^{(k)}(x)$ can be evaluated pointwise because we estimated the normalizations constants of the proposal $\mu^{(k)}_{l^{(k)}}$.
As shown in Appendix B with a martingale argument, because of the optional property of $K_{\mathrm{min}}$, this estimator is unbiased
$\mathbb{E}\,\hat{\gamma}_{l_{\mathrm{max}}}^{H}=\gamma_{l_{\mathrm{max}}}^{\ast},$
and its variance scales in $\mathcal{O}(H^{-2})$. Note that in the present idealized algorithm, samples $X_{k}$’s above are assumed to be drawn exactly from the sequence of importance distributions ${\mu_{l^{(k-1)}}^{(k-1)}}$’s. In practice, as detailed in Section 4, this assumption will not hold exactly because of the adaptive features used in the sampling of the proposals by the Monte Carlo routine AMS, yoelding a small bias of order the inverse of the sample size.
Note that one can also estimate $\gamma^{\ast}_{l}(\varphi)$ for various test function $\varphi$ and in particular estimate the probabilities
$p^{\ast}_{l}\simeq\hat{p}_{l}^{H}\displaystyle\mathop{=}^{{\rm def}}\hat{\gamma}_{l}^{H}(\,{\rm l}{\rm l}),$ (3.2)
for various levels $l$. This can be done with no computational overhead because the true score of each snapshot has already been evaluated during the algorithm. The practical interesting case is when the sequence of target levels $l^{(k)}$ converges almost surely towards the prescribed maximum level of interest $l_{\max}$: $l^{(k)}\xrightarrow[]{k\to+\infty}l_{\max},$ which happens when using the stopping criteria described in Section 2.5 (as long as the reduced score converges to the true score as $k$ grows).

### v3

@@ 2.3 Snapshot sampling and learning function
$X\sim\frac{1}{\mu^{(k)}_{l^{(k)}}(\mathrm{e}^{\tau^{(k)}E^{(k)}})}\mathrm{e}^{\tau^{(k)}E^{(k)}}\mu^{(k)}_{l^{(k)}},$
with $\tau^{(k)}\geq 0$. The above penalization is similar to the so-called learning function in the literature on ’active learning’ [20]. In this work, we simply choose an exponential function of the error $c$ with a penalty parameter $\tau^{(k)}$ adaptively chosen at iteration $k$ of the algorithm. In particular, $\tau^{(k)}$ will be chosen to be high in the beginning of the algorithm and then be reduced to 00 at some point. The idea here is that $\tau^{(k)}$ enables to find a compromise between two objectives: first i) the snapshots are used to refine the reduced model $S^{(k)}$, and ii) the snapshots are also used in estimation by importance sampling. A large $\tau^{(k)}$ is associated with the objective i). Indeed, favoring candidates with high errors in a $\mu^{(k)}_{l^{(k)}}$-distributed sample enables a faster learning of the true score $S^{\ast}$. A small or vanishing $\tau^{(k)}$ is associated with objective ii). Indeed, the distribution $\mu^{(k)}_{l^{(k)}}$ is constructed to be a robust importance sampling distribution for the target $\eta_{l^{(k)}}^{\ast}$ so that states with smaller error are needed for estimation, ensuring a limited variance for importance sampling.
@@
2.4 Bridging two distributions with updated score approximation (optional)
We will now focus on what happens after the update of the reduced model that follows the sampling of a new snapshot. We are given a certain level $l^{(k)}$ and an associated importance distribution $\mu_{l^{(k)}}^{(k)}$. The problem is to find a new initial level which we will denote $l_{b}(k+1)$ and then sample the new importance distribution $\mu_{l_{b}(k+1)}^{(k+1)}$ associated with the updated score approximation $S^{(k+1)}$.
We emphasize that this bridging methodology is only optional. It is perfectly possible (and we will provide some numerical comparisons) to set $l_{b}(k+1)=-\infty$ which amounts to start at the next iteration with an importance distribution $mu_{l_{b}(k+1)}^{(k+1)}=\pi$ given by the reference distribution.
This initial level $l_{b}(k+1)$ will then be used in the subsequent $k+1$ iteration to determine (with definition (2.2)) the critical level at iteration $k+1$: $l^{(k+1)}(c,l_{b}(k+1))$ for $S^{(k+1)}$. This will complete the main loop of the algorithm.
However, if one tries to sample $\mu_{l_{b}(k+1)}^{(k+1)}$ starting from $\mu_{l^{(k)}}^{(k)}$, using some form of importance or sequential sampling, there is an issue with inclusion. For instance, simply setting the initial level ${l_{b}(k+1)}$ to $l^{(k)}$, the inclusion $\{S^{(k+1)}>l^{(k)}\}\subseteq\{S^{(k)}>l^{(k)}\}$ will generally not be valid, making the task infeasible.
@@ 2.4 Bridging two distributions with updated score approximation (optional)
$\mathrm{Ent}\left(\mu_{l_{b}(k+1)}^{(k+1)}\mid\mu_{l^{(k_{b})}}^{(k_{b})}\right)\leq-\ln(1-\theta).$ (2.5)
This condition ensures that a relatively straightforward importance sampling procedure enables to obtain the initial importance sampling distribution $\mu_{l_{b}(k+1)}^{(k+1)}$ of the forthcoming iteration, from a previous importance distribution $\mu_{l^{(k_{b})}}^{(k_{b})}$. In particular, the inclusion condition $\{S^{(k+1)}\geq{l_{b}(k+1)}\}\subseteq\{S^{(k_{b})}>l^{(k_{b})}\}$ will hold true.
In addition of the condition (2.5), we also want to ensure that the two conditions in (2.2) used to define the maximal achievable level $l^{(k+1)}$ hold for this initial level ${l_{b}(k+1)}$. This is necessary to ensure formally that ${l_{b}(k+1)}\leq l^{(k+1)}$. Explicitly:
$\mathrm{Ent}(\check{\eta}^{(k+1)}_{l_{b}(k+1)}\mid\mu^{(k+1)}_{l_{b}(k+1)})\leq c,$ (2.6)
$\mathrm{Ent}(\hat{\eta}^{(k+1)}_{l_{\mathrm{max}}}\mid\mu^{(k+1)}_{l_{b}(k+1)})<+\infty.$ (2.7)
A pair $(k_{b},{l_{b}(k+1)})\in\llbracket 0,k\rrbracket\times\mathbb{R}$ satisfying (2.5), (2.6) and (2.7) determine two consecutive proposal distributions $\mu_{l^{(k_{b})}}^{(k_{b})}$ and $\mu^{(k+1)}_{l_{b}(k+1)}$ “bridging” the score approximations $S^{(k_{b})}$ and $S^{(k+1)}$. This pair can be determined according to various strategies.
  • A trivial feasible pair is to set $k_{b}=0$ such that $l^{(k_{b})}=l_{b}(k+1)=-\infty$. As explained in Section 3, this will account in practice to restart the AMS simulation with the new score $S^{(k+1)}$ from the beginning until it reaches the critical level $l^{(k+1)}(c,-\infty).$ However, this option can lead to a significant computational overhead if the complexity of the reduced score evaluation is not negligible.
  • A non trivial feasible pair of interest is given by the largest (lexicographic) pair $(k_{b},{l_{b}(k+1)})\in\llbracket 0,k\rrbracket\times\mathbb{R}$ satisfying the three conditions (2.5), (2.6) and (2.7). This means first selecting the importance sampling distribution $\mu_{l^{(k_{b})}}^{(k_{b})}$, among the already computed distributions, from the most recent to the oldest, and then to search the largest level ${l_{b}(k+1)}$ such that the three conditions hold. We will see in Section 3 that in practice, this solution will lead to an important computation saving, but will require nevertheless to save the history of the previous AMS simulations.
@@ 2.5 Stopping criteria
In practical situations, a maximum level $l_{\mathrm{max}}$ is setting the range of the computation of interest: one is interested in computing the rare event probability $p^{\ast}_{l_{\mathrm{max}}}$. The level $l^{(k)}(c,l_{b})$ defined in (2.2) is constrained to stay below $l_{\mathrm{max}}$, and the definition of $l^{(k)}(c,l_{b})$ is modified as the minimum between $l_{\mathrm{max}}$ and the critical level given by (2.2). The algorithm can thus continue as long as one wishes, for instance until enough snapshot samples are obtained.
In addition, in order to avoid unnecessarily enriching an already accurate reduced score and obtain computational savings, it is possible to stop updating the reduced score if the worst-case log cost at the critical level $l^{(k)}$ is almost zero and if this critical level is just below $l_{\max}$. More precisely we define the update stopping index111 We call a stopping index any exit criteria of the algorithm that is independent of the future snapshots. as the first index $k$ such that
@@
where $\epsilon>0$ is a given threshold parameter close to the machine precision. This stop-update rule must however be implemented only when using a non-trivial bridging procedure. We will indeed see in Section 3 that using such a stopping of update of the reduced score, the normalization constant is no longer updated resulting in a sustained variance while new snapshots are sampled. Yet this sustained variance always occurs when using a non trivial bridging procedure. We will hence see in Section 3 that using the latter stop-update rule with a non-trivial bridge usually largely compensate additional variance with computational (and memory) savings.
2.6 The estimator
We are now ready to discuss the estimation of the probability $p^{\ast}_{l_{\mathrm{max}}}$. It is more convenient to work with the un-normalized version of the target distribution $\eta^{\ast}_{l_{\mathrm{max}}}$ defined in (1.1), which will be denoted from now on by $\gamma^{\ast}_{l_{\mathrm{max}}}$ and defined by
@@ 2.6 The estimator
for the level of interest ${l_{\mathrm{max}}}\in\mathbb{R}$ and some test function $\varphi$. Note that $p^{\ast}_{l_{\mathrm{max}}}=\gamma^{\ast}_{l_{\mathrm{max}}}(\,{\rm l}{\rm l})$.
The importance sampling estimator $\hat{\gamma}_{l_{\max}}^{H}$ of $\gamma^{\ast}_{l_{\max}}$ is defined using a random iteration index which we will denote by $K_{j_{0}}$. Only the snapshots sampled at and after iteration $K_{j_{0}}+1$ will be used to build our importance sampling estimator. More precisely, $K_{j_{0}}$ must be of optional type, in the sense that $K_{j_{0}}$ is defined as the first iteration index $k$ for which a certain property depending on the snapshots $X_{1},\ldots,X_{k}$ holds true.
We propose the following option: $K_{j_{0}}$ is the first iteration index $k$ such that the score of the past snapshots have reached the level $l_{\max}$ (the rare event) $j_{0}$ times, for some prescribed $j_{0}\geq 1$. At iteration $K_{j_{0}}+H$ (for some $H>0$) of the algorithm, the estimator of $\gamma^{\ast}_{l_{\mathrm{max}}}$ is then given by:
$\hat{\gamma}^{H}_{l_{\mathrm{max}}}\displaystyle\mathop{=}^{{\rm def}}\frac{1}{H}\sum_{k^{\prime}=K_{j_{0}}+1}^{K_{j_{0}}+H}f_{l_{\mathrm{max}}}^{(k^{\prime})}(X_{k^{\prime}+1})\delta_{X_{k^{\prime}+1}},$ (2.9)
where, if we choose for $k^{\prime}\geq K_{j_{0}}$ to set the learning parameter as $\tau=0$, the IS density ratio can be simplified as
$f_{l_{\mathrm{max}}}^{(k^{\prime})}(x)\displaystyle\mathop{=}^{{\rm def}}\,{\rm l}{\rm l}_{S^{\ast}\geq{l_{\mathrm{max}}}}(x)\frac{\mathrm{d}\pi}{d\mu_{l^{(k^{\prime})}}^{(k^{\prime})}}(x).$
Setting a learning parameter $\tau>0$ in the early stages of the algorithm is crucial as the proposal law has a low probability mass over the set of rare events, which implies that snapshots taken so far are still insufficient to learn an accurate reduced model in this rare event regime. However, once $j_{0}$ snapshots have already been sampled from the set of rare events, the choice of $\tau=0$ implies that the snapshot ${X_{k^{\prime}+1}}$ is drawn according to the proposition ${d\mu_{l^{(k^{\prime})}}^{(k^{\prime})}}$, thus favoring the robustness of the IS estimator rather than the accuracy of an already refined reduced model. Note that : i) the entropic criterion in (2.2) and (2.7) ensures that $\eta_{l_{\mathrm{max}}}^{\ast}\ll\mu_{l^{(k)}}^{(k)}$, and ii) $f_{l_{\mathrm{max}}}^{(k)}(x)$ can be evaluated pointwise as long as we have an estimate of the normalizations constant of the proposal $\mu^{(k)}_{l^{(k)}}$.
As shown in Appendix B with a martingale argument, because of the optional property of $K_{j_{0}}$, this estimator is unbiased
$\mathbb{E}\,\hat{\gamma}_{l_{\mathrm{max}}}^{H}=\gamma_{l_{\mathrm{max}}}^{\ast},$
and its variance decreases in $\mathcal{O}(H^{-1})$. Note that samples $X_{k}$ in (2.9) are assumed to be drawn exactly from the sequence of importance distributions ${\mu_{l^{(k-1)}}^{(k-1)}}$. In practice, this assumption will not hold exactly because of the adaptive features used in the sampling of the proposals by the Monte Carlo routine, yielding a small bias of order the inverse of the sample size. This issue is discussed in Section 3.
2.7 The idealized importance sampling algorithm
Based on the concepts introduced in Sections 2.1 to 2.6, we can now describe in this section the algorithm that will lead to an estimation of $\gamma^{\ast}_{l_{\mathrm{max}}}$ based on importance sampling.
Algorithm 1 (Idealized ARMS)
@@ 2.7 The idealized importance sampling algorithm
1:level of interest $l_{\mathrm{max}}$, initial reduced score $S^{(1)}$ and error $E^{(1)}$ functions, budget of snapshots $K$, worst-case log cost threshold $c$, reduced score precision $\epsilon\ll e$, number of hits of rare event before estimation $j_{0}$, quantile $\theta=M/N\in(0,1)$ for the bridge between proposals, boolean $restart$ variable for restarting or bridging after a reduced model update, initial learning parameter $\tau_{0}$triangleright ( initialization)
2:Set $l^{(0)}=l_{b}(1)=-\infty$, $H=0$, $j=0$, $k=0$, $\tau=\tau_{0}$, ${Z_{l}}^{(1)}=1$
@@
5:    Simulate the flow of proposal distributions $\mu^{(k)}_{l}$ for $l\leq l^{(k)}$. The maximal level $l^{(k)}$ must be a level $l\in[l_{b}(k),l_{\mathrm{max}}]$ satisfying both of the following two conditions:
$\mathrm{Ent}(\check{\eta}^{(k)}_{l}\mid\mu^{(k)}_{l})\leq c$
@@
$l^{(k)}$ is defined as the smallest $l\in[l_{b}(k),l_{\mathrm{max}}]$ (in practice $l^{(k)}=l-\delta$ for some $\delta>0$ very small) violating one the above two conditions; if search fails, then $l^{(k)}=l_{\mathrm{max}}$.
@@
15:    if $j<j_{0}$ then update the learning function parameter $\tau$ else set $\tau=0$ end if $\triangleright$ ( importance sampling estimation, see Section 2.6)
16:    if  $j>j_{0}$ then
@@
$\hat{\gamma}^{H}_{l_{\max}}\leftarrow\frac{1}{H}\left((H-1)\hat{\gamma}^{H-1}_{l_{\max}}+Z^{{(k)}}_{l^{(k)}}\,{\rm l}{\rm l}_{S^{\ast}(X)>l_{\max}}\delta_{X}\right)$
@@
22:         if $!restart$ then
23:             Find the largest (lexicographic) pair $(k^{\prime},l^{\prime})\in\llbracket 0,k\rrbracket\times\mathbb{R}$ satisfying the following three conditions:
@@
$\mathrm{Ent}(\check{\eta}^{(k+1)}_{l^{\prime}}\mid\mu^{(k+1)}_{l^{\prime}})\leq c,$
@@
and denote the solution pair by $(k_{b}(k+1),l_{b}(k+1))\leftarrow(k^{\prime},l^{\prime}).$
24:             Simulate the proposal $\mu^{(k+1)}_{l_{b}(k+1)}$ using $\mu^{(k_{b}(k+1))}_{l^{k_{b}(k+1)}}$; then compute the associated normalization $Z^{(k+1)}_{l_{b}(k+1)}$.
25:         else
26:             Set $l_{b}(k+1)=-\infty$
27:             Set $Z^{(k+1)}=1$ and reinitialize $\mu^{(k+1)}_{l_{b}(k+1)}$ with $\pi$
28:         end if
29:    else
30:         Set $l_{b}(k+1)=l^{(k)}$
31:         Set $Z^{(k+1)}=Z^{(k)}$ and $\mu^{(k+1)}_{l_{b}(k+1)}=\mu^{(k)}_{l^{(k)}}$
32:    end if
33:end while
34:Return Estimate $\hat{\gamma}^{H}_{l_{\max}}$
This algorithm is an idealized concept and, in practice, it is not directly computable. Indeed it assumes exact evaluation of expectations and exact sampling according to the proposal and the surrogate distributions which are defined up to a normalization constant.